# Inertial Coordination Games

### v1

@@ A.5  Proposition 3
where the second line follows from $\Phi^{\prime}(\cdot)\leq 1/\sqrt{2\pi}$, and the fifth line follows from $\mu_{t}^{*}<c$ and $\mu_{t+1}^{*}-\mu_{t}^{*}<t^{-1}$. Hence, for any $\epsilon$, there exists $\overline{\sigma}>0$ such that $|\lambda_{t+1}-\lambda_{t}|<\epsilon$ holds for all $t\geq 1$ and $\sigma>\overline{\sigma}$. Note that this argument does not depend on the value of $\theta$ as long as $|\theta|<\infty$ holds.
Appendix B Signals about past play
@@
Appendix C Finite Player Approximation
@@
For any $(\Sigma,\lambda_{0})$ and $\theta\in\mathbb{R}$,
@@
Appendix D Alternative Interpretations of Model
@@
Appendix E Payoff extensions
@@
$u(a_{it},\lambda_{t-1},\theta):=\begin{cases}\alpha\theta+(1-\alpha)\lambda_{t-1}&\text{if }a_{it}=1\\ c&\text{if }a_{it}=0\\ \end{cases}$
with $\alpha\in(0,1).$ The parameter of $\alpha$ can be interpreted as the inverse of the extent of intertemporal strategic complementarity. To simplify the notations, we define the followings:
$\tilde{\theta}(\alpha)$ $:=\frac{\alpha}{1-\alpha}\theta$
$\tilde{c}(\alpha)$ $:=\frac{c}{1-\alpha}$
$\tilde{\sigma}_{t}^{2}(\alpha)$ $:=\left(\frac{\alpha}{1-\alpha}\right)^{2}\sigma_{t}^{2}$
$\tilde{\eta}_{t}^{2}(\alpha)$ $:=\left(\frac{\alpha}{1-\alpha}\right)^{2}\eta_{t}^{2}.$
@@
$RD(\theta):=\mathbbm{1}\left[\theta\geq\frac{c}{\alpha}-\frac{1}{2}\frac{1-\alpha}{\alpha}\right]=\mathbbm{1}[\tilde{\theta}(\alpha)\geq\tilde{c}(\alpha)-(1/2)]$
@@
All the results in the main text hold by replacing $c$, $\sigma_{t}^{2}$, $\eta_{t}^{2}$, and $\bar{\beta}=1+|c-(1/2)-\theta|/2$ with $\tilde{c}(\alpha)$, $\tilde{\sigma}_{t}^{2}(\alpha)$, $\tilde{\eta}_{t}^{2}(\alpha)$, and $\tilde{\beta}:=1+\big{|}\tilde{c}(\alpha)-(1/2)-\tilde{\theta}(\alpha)\big{|}/2$, respectively.
@@
$\mu_{t}^{*}+\Phi\left(\frac{\mu_{t}^{*}-\mu_{t-1}^{*}}{\sqrt{\tilde{\eta}_{t-1}^{2}(\alpha)+\tilde{\eta}_{t}^{2}(\alpha)}}\right)=\tilde{c}(\alpha)$
$\lambda_{t}(\theta)=\mathbbm{1}\left\{\tilde{\theta}(\alpha)\geq\mu_{t}^{*}\right\}=\Phi\left(\frac{\tilde{\theta}(\alpha)-\mu_{t}^{*}}{\tilde{\eta}_{t}(\alpha)}\right)$
At $t=1$, agent $i$’s posterior belief is given by $\theta|I_{i1}\sim N(x_{i1},\sigma_{1}^{2})$, which implies $\tilde{\theta}(\alpha)|I_{i1}\sim N(\tilde{\mu}_{i1},\tilde{\sigma}_{1}^{2}(\alpha))$ with $\tilde{\mu}_{i1}:=\mathbb{E}_{i1}[\tilde{\theta}(\alpha)]=\alpha x_{i1}/(1-\alpha)=:\tilde{x}_{i1}$. Since $\lambda_{0}$ is common knowledge, she takes $a_{i1}=1$ if and only if $\tilde{\mu}_{i1}\geq\tilde{c}-\lambda_{0}=\mu_{1}^{*}$. Since $\tilde{\mu}_{i1}|\tilde{\theta}(\alpha)\sim N(\tilde{\theta}(\alpha),\tilde{\sigma}_{1}^{2}(\alpha))$, $\lambda_{1}$ is given by
$\lambda_{1}=\mathbb{P}_{\theta}(\tilde{\mu}_{i1}\geq\mu_{1}^{*})=1-\Phi\left(\frac{\mu_{1}^{*}-\tilde{\theta}(\alpha)}{\tilde{\sigma}_{1}(\alpha)}\right)=\Phi\left(\frac{\tilde{\theta}(\alpha)-\mu_{1}^{*}}{\tilde{\sigma}_{1}(\alpha)}\right).$
@@
Now, for $t>1$, suppose that $a_{j,t-1}=1$ if and only if $\tilde{\mu}_{j,t-1}\geq\mu_{t-1}^{*}$. By standard Gaussian-Gaussian Bayesian updating, we have $\tilde{\theta}(\alpha)|I_{it}\sim N(\tilde{\mu}_{it},\tilde{\eta}_{t}^{2}(\alpha))$, where
$\tilde{\mu}_{it}$ $=\frac{\tilde{\eta}_{t-1}^{-2}(\alpha)}{\tilde{\eta}_{t-1}^{-2}(\alpha)+\tilde{\sigma}_{t}^{-2}(\alpha)}\tilde{\mu}_{i,t-1}+\frac{\tilde{\sigma}_{t}^{-2}(\alpha)}{\tilde{\eta}_{t-1}^{-2}(\alpha)+\tilde{\sigma}_{t}^{-2}(\alpha)}\tilde{x}_{it}=\tilde{\eta}_{t}^{2}(\alpha)\sum_{s=1}^{t}\tilde{\sigma}_{s}^{-2}(\alpha)\tilde{x}_{is}$
$\tilde{\eta}_{t}^{2}(\alpha)$ $=\frac{\tilde{\eta}_{t-1}^{2}(\alpha)}{1+\tilde{\sigma}_{t}^{-2}(\alpha)\tilde{\eta}_{t-1}^{2}(\alpha)}=\left(\sum_{s=1}^{t}\tilde{\sigma}_{s}^{-2}(\alpha)\right)^{-1},$
where $\tilde{x}_{is}:=\alpha x_{is}/(1-\alpha)$. Then, by an analogous argument as in the proof of Lemma 1, ($\mu^{*}$-LoM) and ($\lambda$-LoM) hold for $t$. Once we obtain ($\mu^{*}$-LoM) and ($\lambda$-LoM), the rest of the arguments carry through by replacing $c$, $\sigma_{t}^{2}$, $\eta_{t}^{2}$, and $\bar{\beta}=1+|c-(1/2)-\theta|/2$ with $\tilde{c}(\alpha)$, $\tilde{\sigma}_{t}^{2}(\alpha)$, $\tilde{\eta}_{t}^{2}(\alpha)$, and $\tilde{\beta}:=1+\big{|}\tilde{c}(\alpha)-(1/2)-\tilde{\theta}(\alpha)\big{|}/2$, respectively. ∎
@@
Theorem 1 holds exactly as stated in the main text, without replacing the above expressions with their $\alpha$-counterparts.
@@
Note that $\eta_{t}^{-2}=O(t^{p})$ if and only if $\tilde{\eta}_{t}^{-2}(\alpha)=O(t^{p})$, and that $\eta_{t}^{-2}=\Omega(t^{q})$ if and only if $\tilde{\eta}_{t}^{-2}(\alpha)=\Omega(t^{q})$. Hence, Theorem 1 holds without replacing the above expressions with their $\alpha$-counterparts. ∎

### v2

@@ A.5  Proposition 3
Hence, for any $\epsilon$, there exists $\overline{\sigma}>0$ such that $|\lambda_{t+1}-\lambda_{t}|<\epsilon$ holds for all $t\geq 1$ and $\sigma>\overline{\sigma}$. Note that this argument does not depend on the value of $\theta$ as long as $|\theta|<\infty$ holds.
Appendix B Small Lags
Consider $\mathcal{T}=\{1,2,\dots\}$ and fix a learning process $\Sigma=(\sigma_{t}^{2})_{t\in\mathcal{T}}$. Now we define an alternative environment with $\mathcal{T}^{(n)}:=\left\{\frac{1}{n},\frac{2}{n},\dots\right\}$ and $\Sigma^{(n)}:=((\sigma_{t}^{(n)})^{2})_{t\in\mathcal{T}^{(n)}}$ where $\sigma_{t}^{(n)}=\sqrt{n}\sigma_{s}$ when $t\in(s-1,s]$. Note that $\sigma_{t}^{(n)}$ is constant in $t$ for all $t\in(s-1,s]$. Then, the posterior variance is given by
$(\eta_{t}^{(n)})^{2}=\left(\sum_{k\in\mathcal{T}^{(n)},k\leq t}\frac{1}{(\sigma_{t}^{(n)})^{2}}\right)^{-1}.$
Using this, we have the following indifference condition and the aggregate time-$t$ play:
$\mu_{t+1/n}^{(n)*}+\Phi\left(\frac{\mu_{t+1/n}^{(n)*}-\mu_{t}^{(n)*}}{\sqrt{(\eta_{t+1/n}^{(n)})^{2}+(\eta_{t}^{(n)})^{2}}}\right)=c,\quad\lambda_{t}^{(n)}(\theta)=\Phi\left(\frac{\theta-\mu_{t}^{(n)*}}{\eta_{t}^{(n)}}\right)$
where $\mu_{t}^{(n)*}$ is time-$t$ threshold belief for $t\in\mathcal{T}^{(n)}$. Define $\lambda^{(n)}_{\infty}(\theta):=\lim_{t\to\infty}\lambda_{t}^{(n)}(\theta)$.
Proposition 4.
The following statements are true:
  1. P1.
    Consider the model with contemporaneous incentives, in which $U(1,t)=\theta+\lambda_{t}$. Then, the unique symmetric threshold strategy is $\mu_{t}^{*}=c-(1/2)$.
  2. P2.
    For each $n\geq 1$ and $t\geq 1$, there exists $\epsilon(n,t)>0$ such that $\mu_{t}^{(n)*}$ is an $\epsilon(n,t)$-equilibrium of the contemporaneous game. Furthermore, for each fixed $t\geq 1$, $\lim_{n\to\infty}\epsilon(n,t)=0$.
  3. P3.
    Fix any $(\lambda_{0},\Sigma,c)$. For any $n\geq 1$, the following hold:
    • (i)
      For all $\epsilon>0$,
[…]
    • (ii)
      For all $\epsilon>0$,
      $\eta^{-2}_{t}=\Omega(t^{2+\epsilon})\implies$ $\lambda_{\infty}^{(n)}(\theta)=NRD(\theta)$
Proof.
We show each part in turn.
Part 1. Fix any $t\in\mathcal{T}$. Suppose that agents follow the same threshold strategy; that is, agent $i$ chooses action $1$ if and only if $\mu_{it}\geq\mu_{t}^{*}$, and this is common knowledge. Agent $i$’s expected payoff of playing $1$ is
$\mathbb{E}_{it}[\theta+\lambda_{t}(\theta)]=\mu_{it}+\mathbb{P}_{it}(\mu_{jt}\geq\mu_{t}^{*})=\mu_{it}+\Phi\left(\frac{\mu_{it}-\mu_{t}^{*}}{\sqrt{2}\eta_{t}}\right).$
Since she is indifferent between playing 00 and $1$ when $\mu_{it}=\mu_{t}^{*}$, it must be that
$\mu_{t}^{*}+\Phi\left(\frac{\mu_{t}^{*}-\mu_{t}^{*}}{\sqrt{2}\eta_{t}}\right)=c\iff\mu_{t}^{*}=c-\frac{1}{2}.$
Part 2. Fix any $n$ and $t\in\mathcal{T}^{(n)}$. Without loss of generality, assume $\lambda_{0}>\frac{1}{2}$ so that $\mu_{t}^{(n)*}\uparrow\mu_{\infty}^{(n)*}$. Suppressing dependencies on these variables, define
$x_{1}(\theta)=\frac{\theta-\mu_{t}^{(n)*}}{\eta_{t}^{(n)}}\quad\text{and}\quad x_{2}(\theta)=\frac{\theta-\mu_{t-1/n}^{(n)*}}{\eta_{t-1/n}^{(n)}}$
and $x(\theta)=\min\Big{\{}|x_{1}(\theta)|,|x_{2}(\theta)|\Big{\}}$. Observe that
$\left|\lambda_{t}^{(n)}(\theta)-\lambda_{t-1/n}^{(n)}(\theta)\right|$ $=\left|\Phi\left(x_{1}(\theta)\right)-\Phi\left(x_{2}(\theta)\right)\right|$
$\leq\varphi(x(\theta))\left|x_{1}(\theta)-x_{2}(\theta)\right|$
$\leq\theta\varphi(x(\theta))\left(\frac{1}{\eta_{t}^{(n)}}-\frac{1}{\eta_{t-1/n}^{(n)}}\right)+\frac{1}{\sqrt{2}\pi}\frac{\eta_{t-1/n}^{(n)}\mu_{t}^{(n)*}-\eta_{t}^{(n)}\mu_{t-1/n}^{(n)*}}{\eta_{t}^{(n)}\eta_{t-1/n}^{(n)}}$ ($\varphi(\cdot)\leq(\sqrt{2}\pi)^{-1}$, $\eta_{t-1/n}^{(n)}\geq\eta_{t}^{(n)}$, and $\mu_{t}^{(n)*}\geq\mu_{t-1/n}^{(n)*}$)
$=\theta\varphi(x(\theta))\left(\frac{1}{\eta_{t}^{(n)}}-\frac{1}{\eta_{t-1/n}^{(n)}}\right)+\frac{\mu_{t}^{(n)*}-\mu_{t-1/n}^{(n)*}}{\sqrt{2}\pi\eta_{t}^{(n)}}+\frac{\eta_{t-1/n}^{(n)}-\eta_{t}^{(n)}}{\sqrt{2}\pi\eta_{t}^{(n)}\eta_{t-1/n}^{(n)}}\mu_{t-1/n}^{(n)*}$
$\leq\theta\varphi(x(\theta))\left(\frac{1}{\eta_{t}^{(n)}}-\frac{1}{\eta_{t-1/n}^{(n)}}\right)+\frac{1}{\sqrt{2}\pi\eta_{t}^{(n)}(nt-1)}$ ($\mu_{t}^{(n)*}-\mu_{t-1/n}^{(n)*}\leq(nt-1)^{-1}$ by (3))
$\leq\max\left\{|\mu_{t}^{(n)*}|+\eta_{t}^{(n)},|\mu_{t-1/n}^{(n)*}|+\eta_{t-1/n}^{(n)}\right\}\left(\frac{1}{\eta_{t}^{(n)}}-\frac{1}{\eta_{t-1/n}^{(n)}}\right)+\frac{1}{\sqrt{2}\pi\eta_{t}^{(n)}(nt-1)}$
$:=\epsilon(n,t),$
where the last inequality follows from
$\theta\varphi(|x_{1}(\theta)|)$ $=\left(\mu_{t}^{(n)*}+\eta_{t}^{(n)}x_{1}(\theta)\right)\varphi(|x_{1}(\theta)|)$
$\leq\left(|\mu_{t}^{(n)*}|+\eta_{t}^{(n)}x_{1}(\theta)\right)\varphi(|x_{1}(\theta)|)$
$\leq|\mu_{t}^{(n)*}|+\eta_{t}^{(n)}$ ($a\varphi(|a|)\leq\varphi(1)\leq 1$)
and $\theta\varphi(|x_{2}(\theta)|)\leq|\mu_{t-1/n}^{(n)*}|+\eta_{t-1/n}^{(n)}$ by analogous arguments.
Hence, the above convergence is uniform in $\theta$; i.e. there exists $\epsilon(n,t)$ with $\epsilon(n,t)\downarrow 0$ as $n\to\infty$ for any fixed $t\geq 1$, such that
$\sup_{\theta\in\mathbb{R}}\left|\lambda_{t}^{(n)}(\theta)-\lambda_{t-1/n}^{(n)}(\theta)\right|\leq\epsilon(n,t)$
Hence, the time-$t$ strategy “$a_{it}=1$ iff $\mu_{it}\geq\mu_{t}^{(n)*}$” is an $\epsilon(n,t)$-equilibrium at time $t$, since we may bound the time-$t$ expected payoff of using this strategy in the contemporaneous game uniformly across players’ time-$t$ information sets $\mu_{it}$: if $\mu_{it}\geq\mu_{t}^{(n)*}$,
$\mathbb{E}_{it}\Big{[}\theta+\lambda_{t}^{(n)}(\theta)\Big{]}=\mathbb{E}_{it}\Big{[}\theta+\lambda_{t-1/n}^{(n)}(\theta)\Big{]}+\mathbb{E}_{it}\Big{[}\lambda_{t}^{(n)}(\theta)-\lambda_{t-1/n}^{(n)}(\theta)\Big{]}$
$\geq c-\mathbb{E}_{it}\Big{[}\lambda_{t-1/n}^{(n)}(\theta)-\lambda_{t}^{(n)}(\theta)\Big{]}$
$\geq c-\mathbb{E}_{it}\Big{|}\lambda_{t-1/n}^{(n)}(\theta)-\lambda_{t}^{(n)}(\theta)\Big{|}$
$\geq c-\epsilon(n,t)\quad\forall\mu_{it}\in\mathbb{R}$
and an exactly analogous argument holds for the case $\mu_{it}<\mu_{t}^{(n)*}$.
Part 3 (i). Without loss of generality, assume $\lambda_{0}>\frac{1}{2}$ so that $\mu_{t}^{(n)*}\uparrow\mu_{\infty}^{(n)*}$. By definition, $\eta^{-2}_{t}=O(t^{2-\epsilon})$ means that there exist $C$ and $T$ such that $\eta_{t}^{-2}\leq Ct^{2-\epsilon}$ for all $t\in\{T,T+1,\dots\}$. Since $(\eta_{t}^{(n)})^{-2}=\eta_{t}^{-2}$ for $t\in\{1,2,\dots\}$, we have $(\eta_{t}^{(n)})^{-2}\leq Ct^{2-\epsilon}$ for all $t\in\{T,T+1,\dots\}$. Hence,
$(\eta_{t}^{(n)})^{2}+(\eta_{t-1/n}^{(n)})^{2}\geq 2(\eta_{t}^{(n)})^{2}\geq 2C^{-1}t^{-(2-\epsilon)}\quad\text{for $t\in\{T,T+1,\dots\}$}$
Furthermore, since $\mu_{t}^{(n)*}-\mu_{t-1/n}^{(n)*}$ is decreasing in $t$, we have
$\mu_{t}^{(n)*}-\mu_{t-1/n}^{(n)*}\leq\frac{1}{nt-1}\sum_{s\in\mathcal{T}^{(n)},s\leq t-1/n}(\mu_{s+1/n}^{(n)*}-\mu_{s}^{(n)*})=\frac{1}{nt-1}(\mu_{t}^{(n)*}-\mu_{1/n}^{(n)*})\leq\frac{1}{nt-1}.$ (3)
Using this inequality, we can apply the squeeze theorem and compute $\mu_{\infty}^{(n)*}$ as follows: for $t\in\{T,T+1,\dots\}$, we have
$0\leq\frac{\mu_{t}^{(n)*}-\mu_{t-1/n}^{(n)*}}{\sqrt{(\eta_{t}^{(n)})^{2}+(\eta_{t-1/n}^{(n)})^{2}}}\leq\underbrace{(C/2)^{1/2}\frac{t^{1-\epsilon/2}}{nt-1}}_{\to 0\quad\text{as $t\to\infty$}}$
$\implies\mu_{\infty}^{(n)*}=c-\Phi\left(\lim_{t\to\infty}\frac{\mu_{t}^{(n)*}-\mu_{t-1/n}^{(n)*}}{\sqrt{(\eta_{t}^{(n)})^{2}+(\eta_{t-1/n}^{(n)})^{2}}}\right)=c-(1/2)$
Note that since the sequence $(\mu_{t}^{(n)*})_{t\in\mathcal{T}^{(n)}}$ converges, the subsequence $(\mu_{t}^{(n)*})_{t\in\mathcal{T}}$ also converges to the same limit $\mu_{\infty}^{(n)*}$.
Part 3 (ii). By definition, $\eta^{-2}_{t}=\Omega(t^{2+\epsilon})$ means that there exist $C$ and $T$ such that $\eta_{t}^{-2}\geq Ct^{2+\epsilon}$ for all $t\in\mathcal{T}$ and $t\geq T$. Since $t^{2+\epsilon}$ is strictly convex, it must be that $(\eta_{t}^{(n)})^{-2}\geq Ct^{2+\epsilon}$ for all $t\in\mathcal{T}^{(n)}$ and $t\geq T$.
To invoke Theorem 1, consider an auxiliary environment in $\mathcal{T}$ such that $\tilde{\eta}_{s}:=\eta_{s/n}^{(n)}$ for $s\in\mathcal{T}$. Then, for all $s\in\mathcal{T}$ and $s\geq nT=:\tilde{T}$, we have
$(\tilde{\eta}_{s})^{-2}$ $=(\eta_{s/n}^{(n)})^{-2}\geq C\left(\frac{s}{n}\right)^{2+\epsilon}$ ($\frac{s}{n}\in\mathcal{T}^{(n)}$ and $\frac{s}{n}\geq T$)
$=\underbrace{Cn^{-(2+\epsilon)}}_{=:\tilde{C}}s^{2+\epsilon},$
which implies $(\tilde{\eta}_{s})^{-2}=\Omega(s^{2+\epsilon})$. Hence, Theorem 1 (ii) implies $\lambda_{\infty}^{(n)}(\theta)=NRD(\theta)$. ∎
Appendix C Signals about past play
@@
ONLINE APPENDIX TO ‘INERTIAL COORDINATION GAMES’
Andrew Koh [^1] Department of Economics, MIT, ajkoh@mit.edu  Ricky Li [^2] Department of Economics, MIT, rickyli@mit.edu
Kei Uzui [^2] Department of Economics, MIT, kuzui@mit.edu
FOR ONLINE PUBLICATION ONLY
Appendix IV Finite Players
@@
For any $N\geq 1$, $(\Sigma,\lambda_{0})$, and $\theta\in\mathbb{R}$,
@@
Appendix V Alternative Interpretations of Model
@@
Appendix VI Payoffs generalizations
@@
$u(a_{it},\lambda_{t-1},\theta):=\begin{cases}a\theta+b\lambda_{t-1}&\text{if }a_{it}=1\\ c&\text{if }a_{it}=0\\ \end{cases}$
with $a\in(0,1)$ and $b\in(0,1)$. The parameter of $b/a$ can be interpreted as the extent of intertemporal strategic complementarity. To simplify the notations, we define the followings:
$\tilde{\theta}(a,b)$ $:=\frac{a}{b}\theta$
$\tilde{c}(a,b)$ $:=\frac{c}{b}$
$\tilde{\sigma}_{t}^{2}(a,b)$ $:=\left(\frac{a}{b}\right)^{2}\sigma_{t}^{2}$
$\tilde{\eta}_{t}^{2}(a,b)$ $:=\left(\frac{a}{b}\right)^{2}\eta_{t}^{2}.$
@@
$RD(\theta):=\mathbbm{1}\left[\theta\geq\frac{c}{a}-\frac{1}{2}\frac{b}{a}\right]=\mathbbm{1}[\tilde{\theta}(a,b)\geq\tilde{c}(a,b)-(1/2)]$
@@
All the results in the main text hold by replacing $c$, $\sigma_{t}^{2}$, $\eta_{t}^{2}$, and $\bar{\beta}=1+|c-(1/2)-\theta|/2$ with $\tilde{c}(a,b)$, $\tilde{\sigma}_{t}^{2}(a,b)$, $\tilde{\eta}_{t}^{2}(a,b)$, and $\tilde{\beta}:=1+\big{|}\tilde{c}(a,b)-(1/2)-\tilde{\theta}(a,b)\big{|}/2$, respectively.
@@
$\mu_{t}^{*}+\Phi\left(\frac{\mu_{t}^{*}-\mu_{t-1}^{*}}{\sqrt{\tilde{\eta}_{t-1}^{2}(a,b)+\tilde{\eta}_{t}^{2}(a,b)}}\right)=\tilde{c}(a,b)$
$\lambda_{t}(\theta)=\mathbbm{1}\left\{\tilde{\theta}(a,b)\geq\mu_{t}^{*}\right\}=\Phi\left(\frac{\tilde{\theta}(a,b)-\mu_{t}^{*}}{\tilde{\eta}_{t}(a,b)}\right)$
At $t=1$, agent $i$’s posterior belief is given by $\theta|I_{i1}\sim N(x_{i1},\sigma_{1}^{2})$, which implies $\tilde{\theta}(a,b)|I_{i1}\sim N(\tilde{\mu}_{i1},\tilde{\sigma}_{1}^{2}(a,b))$ with $\tilde{\mu}_{i1}:=\mathbb{E}_{i1}[\tilde{\theta}(a,b)]=ax_{i1}/b=:\tilde{x}_{i1}$. Since $\lambda_{0}$ is common knowledge, she takes $a_{i1}=1$ if and only if $\tilde{\mu}_{i1}\geq\tilde{c}-\lambda_{0}=\mu_{1}^{*}$. Since $\tilde{\mu}_{i1}|\tilde{\theta}(a,b)\sim N(\tilde{\theta}(a,b),\tilde{\sigma}_{1}^{2}(a,b))$, $\lambda_{1}$ is given by
$\lambda_{1}=\mathbb{P}_{\theta}(\tilde{\mu}_{i1}\geq\mu_{1}^{*})=1-\Phi\left(\frac{\mu_{1}^{*}-\tilde{\theta}(a,b)}{\tilde{\sigma}_{1}(a,b)}\right)=\Phi\left(\frac{\tilde{\theta}(a,b)-\mu_{1}^{*}}{\tilde{\sigma}_{1}(a,b)}\right).$
@@
Now, for $t>1$, suppose that $a_{j,t-1}=1$ if and only if $\tilde{\mu}_{j,t-1}\geq\mu_{t-1}^{*}$. By standard Gaussian-Gaussian Bayesian updating, we have $\tilde{\theta}(a,b)|I_{it}\sim N(\tilde{\mu}_{it},\tilde{\eta}_{t}^{2}(a,b))$, where
$\tilde{\mu}_{it}$ $=\frac{\tilde{\eta}_{t-1}^{-2}(a,b)}{\tilde{\eta}_{t-1}^{-2}(a,b)+\tilde{\sigma}_{t}^{-2}(a,b)}\tilde{\mu}_{i,t-1}+\frac{\tilde{\sigma}_{t}^{-2}(a,b)}{\tilde{\eta}_{t-1}^{-2}(a,b)+\tilde{\sigma}_{t}^{-2}(a,b)}\tilde{x}_{it}=\tilde{\eta}_{t}^{2}(a,b)\sum_{s=1}^{t}\tilde{\sigma}_{s}^{-2}(a,b)\tilde{x}_{is}$
$\tilde{\eta}_{t}^{2}(a,b)$ $=\frac{\tilde{\eta}_{t-1}^{2}(a,b)}{1+\tilde{\sigma}_{t}^{-2}(a,b)\tilde{\eta}_{t-1}^{2}(a,b)}=\left(\sum_{s=1}^{t}\tilde{\sigma}_{s}^{-2}(a,b)\right)^{-1},$
where $\tilde{x}_{is}:=ax_{is}/b$. Then, by an analogous argument as in the proof of Lemma 1, ($\mu^{*}$-LoM) and ($\lambda$-LoM) hold for $t$. Once we obtain ($\mu^{*}$-LoM) and ($\lambda$-LoM), the rest of the arguments carry through by replacing $c$, $\sigma_{t}^{2}$, $\eta_{t}^{2}$, and $\bar{\beta}=1+|c-(1/2)-\theta|/2$ with $\tilde{c}(a,b)$, $\tilde{\sigma}_{t}^{2}(a,b)$, $\tilde{\eta}_{t}^{2}(a,b)$, and $\tilde{\beta}:=1+\big{|}\tilde{c}(a,b)-(1/2)-\tilde{\theta}(a,b)\big{|}/2$, respectively. ∎
@@
Theorem 1 holds exactly as stated in the main text, without replacing the above expressions with their $a,b$-counterparts.
@@
Note that $\eta_{t}^{-2}=O(t^{p})$ if and only if $\tilde{\eta}_{t}^{-2}(a,b)=O(t^{p})$, and that $\eta_{t}^{-2}=\Omega(t^{q})$ if and only if $\tilde{\eta}_{t}^{-2}(a,b)=\Omega(t^{q})$. Hence, Theorem 1 holds without replacing the above expressions with their $a,b$-counterparts. ∎